# Extreme mass ratio inspirals in dark matter halos: Dynamics and distinguishability of halo models

### v1

@@ I Introduction
A particularly promising scenario involves the formation of DM spikes around black holes (BHs). Both Newtonian and relativistic calculations predict that particle accretion can lead to the development of overdensities, which redistribute to form a cusp. The specific characteristics of these cusps depend on the initial DM profile, with their spatial extent determined by the BH mass Gondolo and Silk (1999); Sadeghian et al. (2013).
@@
Modeling EMRIs in vacuum general relativity (GR) is already a formidable challenge, which is still incomplete after three decades of impressive theoretical developmens Barack and Pound (2019). Extending these calculations to include environmental effects is even more complex, primarily due to two key difficulties: (i) the lack of relativistic solutions describing BHs embedded within matter distributions, and (ii) the intricate couplings between metric and matter variables. As a result, most studies investigating environmental effects on EMRIs rely on approximations, such as post-Newtonian methods or geodesic models that neglect dissipative effects Kocsis et al. (2011); Kavanagh et al. (2020); Coogan et al. (2022); Cole et al. (2023); Tomaselli et al. (2023); Berezhiani et al. (2024).
@@ II.1 Background solution
For a given density profile $\rho(r)$, solving Eqs. (5) fully specifies the background spacetime. We follow the numerical approach developed in Ref. Figueiredo et al. (2023) and consider three families of DM distribution, the Hernquist Hernquist (1990), Navarro-Frenk-White (NFW) Navarro et al. (1997) and Einasto Einasto (1965, 1969) models. The density profile of the first two can be described in terms of a two-parameter semianalytic function
@@ II.2.1 Axial perturbations
[…]
@@ III Numerical setup and detectability
$\dot{E}=\sum_{\ell=2}^{\ell_{\text{max}}}\sum_{m=-\ell}^{\ell}(\dot{E}_{\ell m}^{\infty,\text{ax}}+\dot{E}_{\ell m}^{\infty,\text{pol}})\ .$ (30)
The results presented in the main text have been obtained by setting $\ell_{\text{max}}=2$. In Appendix A, we validate this choice by comparing the results for different values of $\ell_{\text{max}}$. The values of $E_{\ell m}^{\rm ax,pol}$ are computed on a uniform grid of points in the range $r\in[6.0001,14.5001]M_{\text{BH}}$, with step size of $r_{\text{step}}=0.1M_{\text{BH}}$. We solve the equations of motion (29) on this grid for a primary mass $M=10^{6}M_{\odot}$ and two values of the secondary mass: $m_{p}=(10,50)M_{\odot}$. We extract fluxes by setting the numerical infinity in Eqs. (15) and (25) to $r_{\infty}=2\times 10^{7}M_{\rm{BH}}$ when $M_{\rm{halo}}=10^{2}M_{\rm{BH}}$, and $r_{\infty}=2\times 10^{8}M_{\rm{BH}}$ when $M_{\rm{halo}}=10^{3}M_{\rm{BH}}$. These values correspond to $2a_{0_{\rm{max}}}$, where $a_{0_{\rm{max}}}$ is the maximum value of $a_{0}$ among the three different choices used to define the compactness for a fixed value of $M_{\rm{halo}}$ (we have also checked that changes in $r_{\infty}$ by one order of magnitude do not significantly affect the GW fluxes). Finally, to ensure a proper comparison with vacuum fluxes (and consequently with GW phases and waveforms), we compute the vacuum fluxes at the same numerical infinity.
@@
A more quantitative and accurate assessment can be made by evaluating waveform changes weighted by the detector sensitivity. To this aim we compute the faithfulness ${\cal F}$ between GW signals from EMRIs in vacuum, and those surrounded by the DM distribution:
@@
[…]
@@
where a dot represents a derivative with respect to the coordinate time, the quadrupole moment tensor is defined as $I_{ij}=2m_{p}/d_{\rm L}z^{i}(t)z^{j}(t)$, $z(t)$ is the geodesic of the secondary expressed in Cartesian spatial coordinates, and $d_{\rm L}$ is the source luminosity distance. We refer the reader to the method section of Maselli et al. (2022) for further technical details on the waveform generation, and on the numerical implementation of Eq. (31). Finally, we fix $f_{\text{min}}=10^{-4}$ Hz in the integral of Eq. (32), while $f_{\text{max}}$ corresponds to the Nyquist frequency of the signal. We assume that two waveform models are distinguishable if the faithfulness is smaller than the critical value ¯$\bar{{\cal F}}\simeq 0.92$, corresponding to the estimated distinguishability threshold for a template depending on ten parameters, observed with a signal-to-noise ratio of $\sim 10$ Chatziioannou et al. (2017).
@@ IV Results
In Fig. 2 (top row) we show the ratio of matter to vacuum fluxes for EMRIs evolving in DM halos with $M_{\rm halo}=10^{2}M_{\odot}$ as a function of the secondary’s orbital radius. Filled (hollow) markers represent systems evolving for one year (six months) in the LISA band before plunge, respectively, with each marker positioned at the initial inspiral radius. Deviations from the Schwarzschild case can reach up to $10^{-3}$, and generally increase (if only mildly) as a function of the initial orbital separation.
@@
In Fig. 3 we repeat the dephasing analysis for halos of $M_{\rm halo}=10^{3}M_{\odot}$. The results are qualitatively and quantitatively similar to those for lighter DM distributions. This confirms that the phase evolution in the halo is primarily driven by its compactness, rather than the total halo mass and size.
@@
While promising, these results do not account for parameter correlations in the waveform model or the impact of LISA noise. To assess the relevance of halo effects in realistic GW observations, we have computed the faithfulness ${\cal F}$ defined in Eq. (31) between waveforms in matter and vacuum for the EMRI configurations analyzed above. Our results are summarized in the matrix diagrams of Figs. 4 and 5, where we list ${\cal F}$ for six months and one year of observation, $M_{\rm halo}=(10^{2},10^{3})M_{\odot}$, and for the three compactness values considered above. For both $m_{p}=10M_{\odot}$ and $m_{p}=50M_{\odot}$, GW signals from binaries evolving in DM environments differ significantly from vacuum, with ${\cal F}\ll\bar{{\cal F}}$ (where ¯$\bar{{\cal F}}=0.92$ is the distinguishability threshold discussed in Sec. III) even in the case of diluted distributions with ${\cal C}=10^{-5}$.
@@
As a final remark, we want to emphasize the importance of using a fully relativistic description to model EMRI dynamics within DM halos. To demonstrate this, we have computed the phase difference between binaries in which the orbital elements in Eqs. (29) evolve with two different schemes: one based on the $\ell=m=2$ component of the GW flux in Eq. (30), and the other using the quadrupolar post-Newtonian (PN) prescription from Ref. Speeney et al. (2022):
@@ V Conclusions
These findings are further supported by a refined analysis we performed to assess the distinguishability of GW signals weighted by the LISA sensitivity curve, in which we computed the faithfulness between waveforms emitted in vacuum and in the presence of DM. We find that six months of observation are already sufficient to confidently discriminate between EMRIs in vacuum and EMRIs evolving in DM profiles. More interestingly, we find that LISA can also differentiate among DM profiles with different radial density distributions.
@@ Acknowledgements.
N.S. and E.B. are supported by NSF Grants No. AST-2307146, PHY-2207502, PHY-090003 and PHY-20043, by NASA Grants No. 20-LPS20-0011 and 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation. A.M. and S.G. acknowledge financial support from MUR PRIN Grants No. 2022-Z9X4XS and No. 2020KB33TP. The Authors acknowledge support from the ITA-USA Science and Technology Cooperation program, supported by the Ministry of Foreign Affairs of Italy (MAECI), grant No. PGR01167. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (https://www.arch.jhu.edu/), which is supported by the NSF Grant No. OAC-1920103.
[…]
profile $[M_{\odot}]$ [year] $\ell_{\rm max}=2$ $\ell_{\rm max}=3$ $\ell_{\rm max}=4$ $\ell_{\rm max}=2$ $\ell_{\rm max}=3$ $\ell_{\rm max}=4$
@@
Table 1: GW dephasing $\Delta\Phi$ and faithfulness ${\cal F}$ with respect to vacuum computed for EMRIs with a primary of $M=10^{6}M_{\odot}$ and a different secondary masses ($m_{p}=10M_{\odot}$ or $m_{p}=50M_{\odot}$), evolving for either one year or six months in different DM profiles. We assume halo masses of $\rm{M}_{\rm{halo}}=10^{2}\rm{M}_{\odot}$ and a compactness of ${\cal C}=10^{-3}$. The dephasing and faithfulness are computed using the GW fluxes in Eq. (30), truncating the sum over $\ell$ at $\ell=\ell_{\rm max}$. The values for $\ell_{\rm max}=2$ correspond to the results discussed in Sec. IV.
@@ Appendix A Dependence on $\ell_{\text{max}}$
In the main text we computed GW fluxes fixing $\ell_{\rm max}$ in Eq. (30) to the leading ($\ell=2$) contribution. In this appendix we assess the impact of higher-order multipoles in the calculations of the dephasing and faithfulness. We compare the results presented in Sec. IV with the values of $\Delta\Phi$ and ${\cal F}$ obtained by including higher multipoles ($\ell_{\mathrm{max}}=2,3,4$) across all three density profiles and for different values of the secondary mass. We focus on EMRIs with halo mass $\mathrm{M}_{\mathrm{halo}}=10^{2}M_{\mathrm{BH}}$ and compactness $\mathcal{C}=10^{-3}$, as these parameters yield the largest deviations from the vacuum scenario. The results are presented in Table 1. While increasing $\ell_{\rm max}$ reduces the dephasing by $\sim 5\%$ relative to the baseline configuration, this effect tends to saturate for all halo models and EMRI configurations. More importantly, the distinguishability between signals in vacuum and matter environments remains unaffected. The faithfulness hardly changes at all when we consider different values of $\ell_{\rm max}$. This confirms the robustness of our main findings.
@@ Appendix B The effect of halo cutoff parameters on the DM detectability
To assess the impact of the NFW cutoff radius and of the choice of $r_{e}$ in Eq. (2) for the Einasto model, we compute the faithfulness with respect to vacuum signals for the same EMRI configurations shown in Fig. 4, varying both $r_{c}$ and $r_{e}$. The first and second rows of Fig. 7 display the values of ${\cal F}$ for a DM halo of $M=100M_{\odot}$, assuming $(r_{c}=a_{0},r_{e}=a_{0})$ and $(r_{c}=5a_{0},r_{e}=a_{0}/2)$, respectively. The values of ${\cal F}$, whether computed against vacuum templates or between different DM models, remain consistent with those in Fig. 4, where we assumed $(r_{c}=5a_{0},r_{e}=a_{0})$. This confirms that the distinguishability of halo signals is largely insensitive to the mass and scale of the distribution, while it primarily depends on the overall halo compactness.

### v2

@@ I Introduction
A particularly promising scenario involves the formation of DM spikes around black holes (BHs). Both Newtonian and relativistic calculations predict that particle accretion can lead to the development of overdensities, which redistribute to form a cusp. The specific characteristics of these cusps depend on the initial DM profile, with their spatial extent determined by the BH mass Gondolo and Silk (1999); Sadeghian et al. (2013); Konoplya and Zhidenko (2022).
@@
Modeling EMRIs in vacuum general relativity (GR) is already a formidable challenge, which is still incomplete after three decades of impressive theoretical developmens Barack and Pound (2019). Extending these calculations to include environmental effects is even more complex, primarily due to two key difficulties: (i) the lack of relativistic solutions describing BHs embedded within matter distributions, and (ii) the intricate couplings between metric and matter variables. As a result, most studies investigating environmental effects on EMRIs rely on approximations, such as post-Newtonian methods or geodesic models that neglect dissipative effects Kocsis et al. (2011); Kavanagh et al. (2020); Coogan et al. (2022); Cole et al. (2023); Tomaselli et al. (2023); Berezhiani et al. (2024); Rahman et al. (2024).
@@ II.1 Background solution
For a given density profile $\rho(r)$, solving Eqs. (5) fully specifies the background spacetime. We follow the numerical approach developed in Ref. Figueiredo et al. (2023) and consider three families of DM distribution Salucci (2019), the Hernquist Hernquist (1990), Navarro-Frenk-White (NFW) Navarro et al. (1997) and Einasto Einasto (1965, 1969) models. The density profile of the first two can be described in terms of a two-parameter semianalytic function
@@ II.2.1 Axial perturbations
$\Psi^{\rm out}_{\ell m}=\Psi_{\ell m}(r_{*}\to\infty,\omega)=\frac{e^{i\omega r_{*}}}{W}\int_{r_{\text{h}}}^{r_{\infty}}\Psi_{\ell m}^{\rm(in)}S^{\rm ax}_{\ell m}dr_{*}\ ,$ (15)
@@ III Numerical setup and detectability
$\dot{E}=\dot{E}^{\infty}+\dot{E}^{H},$ (30)
where $\dot{E}^{\infty}$ and $\dot{E}^{H}$ denote the energy fluxes at infinity and at the horizon, respectively. Here we will neglect the contribution from the horizon fluxes. First of all, including them in our setup would require a different numerical scheme, as their accurate extraction through the shooting method for polar modes becomes challenging near $r\to r_{h}$. Secondly, we have explicitly verified, using direct integration methods for the axial sector, that including the horizon fluxes produces relative dephasings between matter and vacuum evolution smaller than $0.01\%$. Given that the ratio of axial to polar energy fluxes at infinity is similar, we expect the contribution from the polar horizon fluxes to be similarly small, and we conclude that horizon fluxes should be negligible for the configurations we consider. We also note that no matter fluxes are present in our model. The matter distribution is modeled as a stationary, gravitationally bound, and non-dissipative system. Although its internal energy may vary through coupling with gravitational perturbations, this energy is merely redistributed within the fluid. The only genuinely radiative degrees of freedom are those of the gravitational field itself. A net energy flux associated with matter would require dissipative processes such as viscosity or mass shedding, which are not included in our stress-energy tensor. The interaction between the secondary and the fluid may in principle generate local effects, such as drag forces, whose consistent relativistic treatment would require a self-force approach and lies beyond the scope of this work. Hence, we explicitly assume that the binary evolution is driven by a balance law that includes only GW fluxes.
The total flux at infinity is then given by
$\dot{E}^{\infty}=\sum_{\ell=2}^{\ell_{\text{max}}}\sum_{m=-\ell}^{\ell}(\dot{E}_{\ell m}^{\infty,\text{ax}}+\dot{E}_{\ell m}^{\infty,\text{pol}})\ .$ (31)
The results presented in the main text have been obtained by setting $\ell_{\text{max}}=2$. In Appendix A, we validate this choice by comparing the results for different values of $\ell_{\text{max}}$. The values of $\dot{E}_{\ell m}^{\infty}$ are computed on a uniform grid of points in the range $r\in[6.0001,14.5001]M_{\text{BH}}$, with step size of $r_{\text{step}}=0.1M_{\text{BH}}$. We solve the equations of motion (29) on this grid for a primary mass $M=10^{6}M_{\odot}$ and two values of the secondary mass: $m_{p}=(10,50)M_{\odot}$. We extract fluxes by setting the numerical infinity in Eqs. (15) and (25) to $r_{\infty}=2\times 10^{7}M_{\rm{BH}}$ when $M_{\rm{halo}}=10^{2}M_{\rm{BH}}$, and $r_{\infty}=2\times 10^{8}M_{\rm{BH}}$ when $M_{\rm{halo}}=10^{3}M_{\rm{BH}}$. These values correspond to $2a_{0_{\rm{max}}}$, where $a_{0_{\rm{max}}}$ is the maximum value of $a_{0}$ among the three different choices used to define the compactness for a fixed value of $M_{\rm{halo}}$ (we have also checked that changes in $r_{\infty}$ by one order of magnitude do not significantly affect the GW fluxes). Finally, to ensure a proper comparison with vacuum fluxes (and consequently with GW phases and waveforms), we compute the vacuum fluxes at the same numerical infinity.
@@
[…]
@@
${\langle h_{m}|h_{v}\rangle}=4\Re\left[\int_{f_{\text{min}}}^{f_{\text{max}}}\frac{\tilde{h}_{m}(f)\tilde{h}_{v}^{*}(f)}{S_{n}(f)}df\right]\,.$ (33)
@@
where a dot represents a derivative with respect to the coordinate time, the quadrupole moment tensor is defined as $I_{ij}=2m_{p}/d_{\rm L}z^{i}(t)z^{j}(t)$, $z(t)$ is the geodesic of the secondary expressed in Cartesian spatial coordinates, and $d_{\rm L}$ is the source luminosity distance. We refer the reader to the method section of Maselli et al. (2022) for further technical details on the waveform generation, and on the numerical implementation of Eq. (32). Finally, we fix $f_{\text{min}}=10^{-4}$ Hz in the integral of Eq. (33), while $f_{\text{max}}$ corresponds to the Nyquist frequency of the signal. We assume that two waveform models are distinguishable if the faithfulness is smaller than the critical value $\bar{{\cal F}}\simeq 0.95$, corresponding to the estimated distinguishability threshold for a template depending on ten parameters, observed with a signal-to-noise ratio of $\sim 10$ Chatziioannou et al. (2017).
@@ IV Results
In Fig. 2 (top row) we show the ratio of matter to vacuum fluxes for EMRIs evolving in DM halos with $M_{\rm halo}=10^{2}M_{\text{BH}}$ as a function of the secondary’s orbital radius. Filled (hollow) markers represent systems evolving for one year (six months) in the LISA band before plunge, respectively, with each marker positioned at the initial inspiral radius. Deviations from the Schwarzschild case can reach up to $10^{-3}$, and generally increase (if only mildly) as a function of the initial orbital separation.
@@
In Fig. 3 we repeat the dephasing analysis for halos of $M_{\rm halo}=10^{3}M_{\text{BH}}$ . The results are qualitatively and quantitatively similar to those for lighter DM distributions. This confirms that the phase evolution in the halo is primarily driven by its compactness, rather than the total halo mass and size.
@@
To assess the impact of the redshift correction, we computed redshifted frequencies [^1] Since the binary orbital frequency is related to the radius by $\Omega=r^{-3/2}$ the radius also transforms as $r_{r}=\frac{r_{v}}{(1+\delta z)^{2/3}}$. and vacuum fluxes following the prescription of Ref. Cardoso et al. (2022b), such that $\dot{E}^{\infty}_{r}\rightarrow(1-\delta\mathcal{C})\,\dot{E}_{v}$, and $\Omega_{r}\rightarrow\Omega_{v}(1+\delta z)$. The numerical factor $\delta$ depends on the DM density profile; we adopted values of $\delta=1,0.9,3$ for the Hernquist, NFW, and Einasto profiles, respectively, following the results of Ref. Figueiredo et al. (2023) for the axial sector. For polar perturbations, obtaining analytical estimates of $\delta$ is more challenging. However, we have verified that varying $\delta$ around these fiducial values does not qualitatively change our results.
As expected, the dephasing decreases when the redshift correction is included (see Table 1 ), but it remains significant. This indicates that, while couplings with matter in the emitted fluxes are local, their secular effects influence the long-term evolution of the binary.
$C=10^{-3}$ $C=10^{-4}$ $C=10^{-5}$
density $m_{p}$ T $\Delta\Phi=$ $\Delta\Phi=$ $\Delta\Phi=$
profile $M_{\odot}] [$year] $\Phi_{m}-\Phi_{v}(\Phi_{r})$ $\Phi_{m}-\Phi_{v}(\Phi_{r})$ $\Phi_{m}-\Phi_{v}(\Phi_{r})$
Hernquist 10 1 436 (299) 89 (75) 54 (52)
0.5 256 (189) 50 (44) 30 (29)
50 1 269 (149) 52 (40) 30 (29)
0.5 168 (103) 34 (27) 21 (20)
NFW 10 1 386 (263) 84 (72) 54 (52)
0.5 226 (130) 48 (42) 30 (29)
50 1 237 (166) 49 (38) 30 (29)
0.5 149 (90) 32 (26) 20 (20)
Einasto 10 1 1360 (953) 182 (141) 64 (60)
0.5 803 (604) 106 (74) 36 (34)
50 1 846 (489) 110 (86) 36 (33)
0.5 526 (331) 70 (50) 24 (22)
Table 1: GW dephasing $\Delta\Phi$ with respect to vacuum (redshifted vacuum) computed for EMRIs with a primary of $M=10^{6}M_{\odot}$ and different secondary masses ($m_{p}=10M_{\odot}$ or $m_{p}=50M_{\odot}$), evolving for either one year or six months in different DM profiles. We assume halo masses of $\rm{M}_{\rm{halo}}=10^{2}\rm{M}_{\odot}$.
[…]
@@
Finally, we verified that all results obtained from the faithfulness analysis remain unchanged when using redshifted fluxes, as done for the dephasing. This confirms that the distinguishability of the signals persists even when the redshift effect is accounted for.
As a final remark, we want to emphasize the importance of using a fully relativistic description to model EMRI dynamics within DM halos. To demonstrate this, we have computed the phase difference between binaries in which the orbital elements in Eqs. (29) evolve with two different schemes: one, $\Phi_{\text{R}}$, based on the $\ell=m=2$ component of the GW flux in Eq. (31), and the other, $\Phi_{\text{PN}}$, using the quadrupolar post-Newtonian (PN) prescription from Ref. Speeney et al. (2022):
@@ V Conclusions
These findings are further supported by a refined analysis we performed to assess the distinguishability of GW signals weighted by the LISA sensitivity curve, in which we computed the faithfulness between waveforms emitted in vacuum and in the presence of DM. We find that six months of observation are already sufficient to confidently discriminate between EMRIs in vacuum and EMRIs evolving in DM profiles. More interestingly, we find that LISA can also differentiate among DM profiles with different radial density distributions. Furthermore, by including the redshift correction, we have verified that the associated changes in the orbital frequency and energy flux tend to reduce the dephasing; however, they do not alter the qualitative behavior of the signal. This confirms that the cumulative impact of local matter couplings over time remains significant for the long-term orbital evolution.
@@ Acknowledgements.
We would thank the anonymous Referee for their valuable comments which have improved the quality of our manuscript. N.S. and E.B. are supported by NSF Grants No. AST-2307146, PHY-2207502, PHY-090003 and PHY-20043, by NASA Grants No. 20-LPS20-0011 and 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation. A.M. and S.G. acknowledge financial support from MUR PRIN Grants No. 2022-Z9X4XS and No. 2020KB33TP. The Authors acknowledge support from the ITA-USA Science and Technology Cooperation program, supported by the Ministry of Foreign Affairs of Italy (MAECI), grant No. PGR01167. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (https://www.arch.jhu.edu/), which is supported by the NSF Grant No. OAC-1920103.
density $m_{p}$ T $\Delta\phi$ $\Delta\phi$ $\Delta\phi$ $\mathcal{F}$ $\mathcal{F}$ $\mathcal{F}$
profile $M_{\odot}] [$year] $\ell_{\rm max}=2$ $\ell_{\rm max}=3$ $\ell_{\rm max}=4$ $\ell_{\rm max}=2$ $\ell_{\rm max}=3$ $\ell_{\rm max}=4$
@@
[…]
@@ Appendix A Dependence on $\ell_{\text{max}}$
[…]
@@ Appendix B The effect of halo cutoff parameters on the DM detectability
[…]